Two wires, each of length 1.2 m, are stretched between two fixed supports. On wire A there is a second-harmonic standing wave whose frequency is 731 Hz. However, the same frequency of 731 Hz is the third harmonic on wire B. Find the speed at which the individual waves travel on each wire.
Answer #1

1:- In case of wire A
the second harmonic frequency is 731 hz, so the fundamental frequency is 365.5.
wavelength,n = 1.2 m
So, velocity,v = f * n = 365.5 * 1.2 = 438 m/sec

2:- In case of Wire B
the second harmonic frequency is 731 hz, so the fundamental frequency is 243.67.
n = 2 *L / 3 = 0.8 m

v = 243.67*0.8 = 194.94 m/sec